# CraftGraffiti: Exploring Human Identity with Custom Graffiti Art via Facial-Preserving Diffusion Models

### v1

@@ 1 Introduction
Human identity preservation in the GenAI era: bias, aesthetics, and open experimentation: Generative systems such as GANs and diffusion models replicate and amplify cultural biases from their training data, entering into the risk of distorting human identity across multiple dimensions (ghosh2024don, ). Since usually trained on not particularly curated datasets, it is known that current generative systems usually tend to disproportionately produce light-skinned, youthful, conventionally attractive faces, marginalizing other demographics, and concentrating on a narrow band of a specific concept of aesthetics (munoz2023uncovering, ). On the other hand, both children and older adults are usually either underrepresented or not properly identified, even for age-progression models (munoz2023uncovering, ; li2018glca, ), presenting results that tend to transform the image of a young person into an adult face, or conversely, elderly people as unrealistically younger figures. Errors in the gender identification are usually more prone to affect women than men, and reinforce gender norms and heteronormativity, as seen in occupational and presentational stereotypes (sun2023smiling, ; zhou2024bias, ). Rooted in imbalanced datasets, these distortions risk homogenizing outputs unless addressed through curated data, fairness-aware training, and interdisciplinary critique (ferrara2023fairness, ).
@@
For this reason, and in light of the relevant risks identified, it is essential to openly address the tensions to which our human identity is exposed in the new digital era; an era of interconnected individuals endowed with generative capacities. Such a topic is profound and complex from the philosophical, ethical, political, and technical dimensions (binns2017fairness, ; alkfairy2024ethical, ). We state that in order to address the discussion and public debate of these topics, we can make use of two main instruments: 1) An artistic installation providing a representation of the customized human pose, technically sound and preserving facial attributes as much as possible, and 2) an ecological environment that allows for the natural interaction with people, providing the explicit context of a living lab or open experimentation space.
@@
CraftGraffiti - Integrating Cultural Expression and Facial Consistency in Generative Street Art: For the first aim, we need a cultural bind that serves as an entry point for people to interact with this new era of GenAI, which has to be inclusive and accessible for all; from this perspective, graffiti appears as an excellent candidate. Graffiti is a culturally significant art form, merging creative expression with the diversity of human identity (das2023power, ; forster2012evaluating, ) and often serving as a voice for marginalized communities (bates2014bombing, ). While recent advances in GANs (xu2021drb, ; yang2019controllable, ) and diffusion models (zhang2023inversion, ; wang2023stylediffusion, ) offer automated generation, existing methods struggle to preserve facial attributes across varying styles and poses. Text-based image editing (kawar2023imagic, ; huang2025diffusion, ) and style transfer models (huang2024diffstyler, ; wang2023stylediffusion, ; guo2024focus, ) fail to simultaneously adjust poses, while pose editing approaches (cheong2023upgpt, ; okuyama2024diffbody, ) lack unified style transfer. Unified models (pascual2024enhancing, ; mokady2023null, ; dong2023prompt, ) attempt both but often lose facial details due to self-attention limitations.
@@
We introduce CraftGraffiti, an end-to-end generative framework that produces custom graffiti-style portraits of human characters while explicitly preserving the subject’s facial identity (see Fig.˜1). This system is text-guided, allowing a user to specify high-level context (e.g., prompting a portrait of the person as a DJ, guitarist, or singer) and to customize the character’s pose, all within a graffiti art style. To accomplish this, CraftGraffiti integrates several components: a CLIP-based text encoder for semantic guidance, a graffiti-style fusion module implemented via a pre-trained diffusion transformer  [^1] https://civitai.com/models/1058970/graffiti-style-fluxFaces fine-tuned on graffiti aesthetics, a LoRA-based adapter (hu2022lora, ) that efficiently injects pose information without requiring full model retraining, and a novel face-consistent self-attention mechanism that ensures key facial features of the input are maintained throughout the diffusion process. The final output is rendered by a variational autoencoder (VAE) decoder, yielding a high-quality graffiti portrait that retains the individual’s identity. This architecture enables simultaneous pose adjustment and style transfer while keeping the person’s face recognizable. Our contributions include: (1) a unified framework for pose-guided, graffiti-themed portrait generation that preserves individual identity; (2) a novel self-attention module specifically designed for facial consistency in diffusion-based generation; and (3) state-of-the-art performance in both qualitative visual fidelity and quantitative identity-preservation metrics compared to existing methods.
@@
A public artistic installation in an ecological environment as a living lab: For our second aim, we propose to explore ecological environments to enhance user-experience experimentation by enabling real-world engagement in authentic contexts (kieffer2017ecoval, ). Living labs provide user-centred, open innovation ecosystems with real-world experimentation, co-creation, and multi-stakeholder collaboration (compagnucci2021livinglabs, ; livinglab_wikipedia, ). Outdoor labs with pervasive sensing and crowdsensing further increase scalability and context-awareness (pournaras2021crowdsensing, ), fostering inclusivity, diversity, and richer insights than artificial environments. From this perspective, music festivals appear as excellent ecological experimentation spaces for this type of trans- and multi-disciplinary investigation. In our case, we set our installation during the four days of the Festival Cruïlla, taking place 9-12 July 2025 (Please refer to the appendix for more details of the installation setup, outcomes, and human feedback).
@@ 2 Preliminaries
Diffusion Models: Text-to-image diffusion models, such as Stable Diffusion rombach2022high ; podell2023sdxl ; everaert2023diffusion , generate images by learning a denoising process that maps latent noise to realistic images ($\mathbf{I}$), conditioned on textual input. Let $\mathbf{x}_{0}\in\mathbb{R}^{H\times W\times C}$ denote a clean image and $\mathbf{x}_{t}$ its noisy version at timestep $t\in\{1,\dots,T\}$. The forward diffusion process ($q(\cdot)$) adds Gaussian noise ($\mathcal{N}(\cdot)$) in a Markov chain:
@@
where $\mathbf{c}$ is a conditioning vector, often a CLIP-encoded text prompt. Training minimizes the error between predicted and true noise or the original image $\mathbf{x}_{0}$, depending on parameterization (mean: $\boldsymbol{\mu}_{\theta}$, and covariance $\boldsymbol{\Sigma}_{\theta}$). Latent diffusion models (LDMs) (wu2023latent, ; nam2024contrastive, ; zhang2023sine, ) perform this process in a compressed latent space $\mathcal{Z}$, using an encoder $\mathcal{E}:\mathbb{R}^{H\times W\times C}\rightarrow\mathcal{Z}$ and decoder $\mathcal{D}:\mathcal{Z}\rightarrow\mathbb{R}^{H\times W\times C}$, improving computational efficiency while preserving generation quality.
Style Transfer: Style transfer (wang2023interactive, ; hamazaspyan2023diffusion, ; everaert2023diffusion, ) aims to modify the appearance of a generated image $\mathbf{x}$ to match the style of a reference image $\mathbf{s}$, while maintaining the content from a source image $\mathbf{c}$. In feature space $\phi(\cdot)$, the style and content losses are defined as:
@@
where $G^{(l)}$ denotes the Gram matrix li2017universal of features at layer $l$, capturing second-order statistics. The total loss combines both objectives:
@@
Pose Customization with LoRA: Generation of characterization conditions on a specified human pose $\mathbf{p}$, often represented as a 2D keypoint skeleton or heatmap (wang2024stable, ; ju2023humansd, ; khandelwal2024reposedm, ). This signal can be embedded and injected into a diffusion model via attention layers or early-stage concatenation. LoRA (hu2022lora, ) offers an efficient means of adapting large diffusion models to new tasks (e.g., pose-guided generation) by injecting low-rank updates into pretrained weight matrices. For a weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA introduces trainable low-rank matrices $A\in\mathbb{R}^{d\times r}$, $B\in\mathbb{R}^{r\times k}$, modifying the weights as:
@@
where $\alpha$ is a scaling factor and $r\ll\min(d,k)$. By freezing base weights and only training $A$ and $B$, LoRA allows efficient fine-tuning with significantly fewer parameters. LoRA can be applied to attention matrices $W_{q},W_{k},W_{v}$, enabling fast pose-driven customization without full model retraining. In CraftGraffiti instead of using keypoints or heatmap, we extend the input prompt via T5 (narang2021transformer, ) and get the cross-attention map as a guided signal ($B$) for the pose (see Fig.˜2). We maintain the identity of the character through the face-consistent self-attention ($A$) and update the self and cross attention of the diffusion transformer via fine-tuning with LoRA for multi-scale latent feature generation. This multi-scale latent feature is further denoised via VAE in order to get the final edited image in graffiti style with an accurate musical pose described in the text prompt.
@@ 3.1 Why should we perform graffiti style transfer before facial attribute preservation
Theorem. Let $I$ be the input face with attributes $a=F_{a}(I)$. Let $S$ be a diffusion-style operator and $P$ an operator satisfying $F_{a}(P(X))=a$ for all $X$ (so $P$ restores $I$’s attributes). Then the image $X^{\prime}=P(S(I))$ obeys $F_{a}(X^{\prime})=a$ (it exactly preserves the original attributes), whereas $X^{\prime\prime}=S(P(I))$ satisfies $F_{a}(X^{\prime\prime})=F_{a}(S(I))$ and in general $F_{a}(S(I))\neq a$ unless $S$ itself preserved $F_{a}$. Hence $(P\circ S)(I)$ maintains the facial attributes exactly, whereas $(S\circ P)(I)$ need not.
@@ 3.2 How the extra dimension in the self-attention helps in facial attribute preservation
In Diffusion models (zhou2024storydiffusion, ; wang2024autostory, ), each self-attention layer computes attention only among features of the current latent (no cross-latent computation). Concretely, the latent feature map of size $H\times W$ is reshaped into a sequence of tokens, and for each token, a query ($Q$), key ($K$), and value ($V$) vector are computed from the same latent features. The attention weights are then obtained by taking dot products of $Q$ and $K$ across all positions (scaled by affinity distance $\sqrt{d}$) and passing through softmax, and the result reweights the $V$ vectors (see Fig.˜3). This means every pixel in the latent try to see every other pixel in the same latent when computing self-attention.
@@ 4 Experimental Results
Implementation Details: We implement CraftGraffiti in PyTorch paszke2019pytorch , utilizing a pre-trained .1 dev  [^3] https://huggingface.co/spaces/black-forest-labs/FLUX.1-dev (12 billion parameter) transformer specifically designed for graffiti style transfer. The guidance scale is set to 7.5. For preserving facial attributes while fine-tuning the transformer with low rank adaptation, the subject guidance factor ($\lambda$) is set to 0.95, and the style intensity factor is set to 0.7. We set the rank ($\sigma$) to 64 and the regularization factor ($\alpha$) to 128 while fine-tuning the transformer for pose customization. We run the denoising process with 100 steps by default. We only perform latent composition in the first quarter of the denoising process (the first 25 steps). All experiments ran on a single NVIDIA A6000 48GB GPU.
Dataset and Evaluation Metric: For graffiti style generation, we have used the images of the 17K-Graffiti dataset  [^4] https://github.com/visual-ds/17K-Graffiti to fine-tune our style fusion diffusion with LoRA and use it as a pre-trained model during inference. Also, to further validate CraftGraffiti, we have gathered images of people from our laboratory, the Computer Vision Center, Barcelona  [^5] https://www.cvc.uab.es/, with their consent, and use them in the paper for validation and pose customization. Furthermore, to perform a fair qualitative analysis, we test CraftGraffiti and other state-of-the-art approaches with facial feature consistency (FFC) (implementation in the Appendix), aesthetic score (Aes) (murray2012ava, ) calculated using LAION aesthetic classifier, and Human Preference Score (HPS) (wu2023human, ). Also, we have compared them against the inference time to understand their real-time use case scenario.
Qualitative Analysis: We compare CraftGraffiti against several FLUX-based approaches, FLUX + IP-Adapter (ye2023ip, ), and FLUX.1 Kontext (labs2025flux, ), however, both approaches fail to generate the graffiti style. Although FLUX + IP-Adapter (ye2023ip, ) tried to maintain facial consistency, it was unable to change style. On the other hand, FLUX.1 Kontext (labs2025flux, ) tries to change the background into a more cartoonist style, and the faces are not consistent. This shows how hard to unify the style transfer and consistency tasks and justify our decision to perform style transfer before facial consistency (see Fig.˜4).
@@
We also compare CraftGraffiti with InstructPix2Pix (brooks2023instructpix2pix, ); however, it neither adds objects nor generates high-quality images. Similarly, VLMs (Grok 3 (grok3beta, ) and GPT4o (hurst2024gpt, )) maintain consistency, unable to blend style due to the complexity of graffiti style transfer. This empirically shows the complexity of the problem (facial consistency in graffiti style) we are tackling and demonstrates the correctness of the component choice of CraftGraffitii for accurate style blending by preserving the facial features (For more qualitative examples, please refer to the Fig.˜10 of the appendix.)
Quantitative Analysis: Table˜1 presents a quantitative comparison against the state-of-the-art image editing techniques. CraftGraffiti stands in the fourth position after GPT4 (hurst2024gpt, ), GPT4o (hurst2024gpt, ), and (ye2023ip, ) respectively, in terms of facial feature consistency (FFC), as it is very difficult to compare the preservation of facial features, as most of the feature extraction models (InceptionResNetV1) consider global features rather than the local ones. However, during style transfer, it is hard to preserve the global features of the input images. As other image editing (grok3beta, ; hurst2024gpt, ; ye2023ip, ) models are unable to blend the graffiti style, preserve the global features, and perform better in this metric.
@@
FLUX + IP-Adapter (ye2023ip, ) 0.8324 3.2414 0.3012 8.2
FLUX.1 Kontext (labs2025flux, ) 0.6741 2.1749 0.2911 6.7
InstructPix2Pix (brooks2023instructpix2pix, ) 0.7112 2.7272 0.1918 3.1
GPT4o (hurst2024gpt, ) 0.8761 4.5193 0.3412 13.4
Grok 3 (grok3beta, ) 0.8513 4.1652 0.3102 11.2
@@
On the other hand, Aesthetic Score (murray2012ava, ; banerjee2025svgcraft, ) and Human Preference Score (wu2023human, ) measure how perfectly the models blend the graffiti style with the input images. CraftGraffiti outperforms all the state-of-the-art image editing techniques in both metrics, as they are unable to blend the graffiti style properly over the input images. Last but not least, CraftGraffiti has very little inference time compared to the VLMs (grok3beta, ; hurst2024gpt, ), which makes it suitable for real-time applications.
@@
Similarly, we perform an ablation study on the self-attention (see Fig.˜6) to preserve the facial attribute during pose customization to understand why the extension of the attention dimension helps to preserve the facial features. In order to do that, we took the generated images with CraftGraffiti and pass them through the self-attention layer of FLUX.1 dev, Subject-driven self-attention layer of Consistory (tewel2024training, ), and our face consistent self-attention layer. It has been observed that the traditional self-attention of FLUX.1 dev focuses on a certain point, whereas Subject-driven Self-Attention of ConsiStory (tewel2024training, ) attends the global features rather than the local ones. On the other hand, our proposed face-consistent self-attention, focused on the local features of the facial attributes, helps to maintain facial attributes during face customization.
Human Evaluation: We have conducted the perceptual study in order to evaluate the credibility and widespread adoption of the graffiti poster synthesized with CraftGraffiti. We chose 60 graffiti posters (20 for each pose) synthesized by (ye2023ip, ; labs2025flux, ; brooks2023instructpix2pix, ; grok3beta, ; hurst2024gpt, ) and CraftGraffiti and asked the users to rate the graffiti on a scale of 1 to 5 (1 Low; 5 high) based on aesthetics, style blending, and recognizability. 47 anonymous users participated in this study. The outcome of this user study is reported in Fig.˜7.

### v2

@@ 1 Introduction
Human identity preservation in the GenAI era: bias, aesthetics, and open experimentation: Generative systems such as GANs and diffusion models replicate and amplify cultural biases from their training data, entering into the risk of distorting human identity across multiple dimensions (Ghosh et al., 2024). Since usually trained on not particularly curated datasets, it is known that current generative systems usually tend to disproportionately produce light-skinned, youthful, conventionally attractive faces, marginalizing other demographics, and concentrating on a narrow band of a specific concept of aesthetics (Muñoz et al., 2023). On the other hand, both children and older adults are usually either underrepresented or not properly identified, even for age-progression models (Muñoz et al., 2023; Li et al., 2018), presenting results that tend to transform the image of a young person into an adult face, or conversely, elderly people as unrealistically younger figures. Errors in the gender identification are usually more prone to affect women than men, and reinforce gender norms and heteronormativity, as seen in occupational and presentational stereotypes (Sun et al., 2023; Zhou, 2024). Rooted in imbalanced datasets, these distortions risk homogenizing outputs unless addressed through curated data, fairness-aware training, and interdisciplinary critique (Ferrara, 2023).
@@
For this reason, and in light of the relevant risks identified, it is essential to openly address the tensions to which our human identity is exposed in the new digital era; an era of interconnected individuals endowed with generative capacities. Such a topic is profound and complex from the philosophical, ethical, political, and technical dimensions (Binns, 2017; Al-kfairy et al., 2024). We state that in order to address the discussion and public debate of these topics, we can make use of two main instruments: 1) An artistic installation providing a representation of the customized human pose, technically sound and preserving facial attributes as much as possible, and 2) an ecological environment that allows for the natural interaction with people, providing the explicit context of a living lab or open experimentation space.
@@
CraftGraffiti - Integrating Cultural Expression and Facial Consistency in Generative Street Art: For the first aim, we need a cultural bind that serves as an entry point for people to interact with this new era of GenAI, which has to be inclusive and accessible for all; from this perspective, graffiti appears as an excellent candidate. Graffiti is a culturally significant art form, merging creative expression with the diversity of human identity (Das, 2023; Forster et al., 2012) and often serving as a voice for marginalized communities (Bates, 2014). While recent advances in GANs (Xu et al., 2021; Yang et al., 2019) and diffusion models (Zhang et al., 2023a; Wang et al., 2023b) offer automated generation, existing methods struggle to preserve facial attributes across varying styles and poses. Text-based image editing (Kawar et al., 2023; Huang et al., 2025) and style transfer models (Huang et al., 2024; Wang et al., 2023b; Guo and Lin, 2024) fail to simultaneously adjust poses, while pose editing approaches (Cheong et al., 2023; Okuyama et al., 2024) lack unified style transfer. Unified models (Pascual et al., 2024; Mokady et al., 2023; Dong et al., 2023) attempt both but often lose facial details due to self-attention limitations.
@@
We introduce CraftGraffiti, an end-to-end generative framework that produces custom graffiti-style portraits of human characters while explicitly preserving the subject’s facial identity (see Fig.˜1). This system is text-guided, allowing a user to specify high-level context (e.g., prompting a portrait of the person as a DJ, guitarist, or singer) and to customize the character’s pose, all within a graffiti art style. To accomplish this, CraftGraffiti integrates several components: a CLIP-based text encoder for semantic guidance, a graffiti-style fusion module implemented via a pre-trained diffusion transformer  [^1] https://civitai.com/models/1058970/graffiti-style-fluxFaces fine-tuned on graffiti aesthetics, a LoRA-based adapter (Hu et al., 2022) that efficiently injects pose information without requiring full model retraining, and a novel face-consistent self-attention mechanism that ensures key facial features of the input are maintained throughout the diffusion process. The final output is rendered by a variational autoencoder (VAE) decoder, yielding a high-quality graffiti portrait that retains the individual’s identity. This architecture enables simultaneous pose adjustment and style transfer while keeping the person’s face recognizable. Our contributions include: (1) a unified framework for pose-guided, graffiti-themed portrait generation that preserves individual identity; (2) a novel self-attention module specifically designed for facial consistency in diffusion-based generation; and (3) state-of-the-art performance in both qualitative visual fidelity and quantitative identity-preservation metrics compared to existing methods.
@@
A public artistic installation in an ecological environment as a living lab: For our second aim, we propose to explore ecological environments to enhance user-experience experimentation by enabling real-world engagement in authentic contexts (Kieffer, 2017). Living labs provide user-centred, open innovation ecosystems with real-world experimentation, co-creation, and multi-stakeholder collaboration (8; 29). Outdoor labs with pervasive sensing and crowdsensing further increase scalability and context-awareness (Pournaras et al., 2021), fostering inclusivity, diversity, and richer insights than artificial environments. From this perspective, music festivals appear as excellent ecological experimentation spaces for this type of trans- and multi-disciplinary investigation. In our case, we set our installation during the four days of the Festival Cruïlla, taking place 9-12 July 2025 (Please refer to the appendix for more details of the installation setup, outcomes, and human feedback).
@@ 2 Preliminaries
Diffusion Models: Text-to-image diffusion models, such as Stable Diffusion Rombach et al. (2022); Podell et al. (2023); Everaert et al. (2023), generate images by learning a denoising process that maps latent noise to realistic images ($\mathbf{I}$), conditioned on textual input. Let $\mathbf{x}_{0}\in\mathbb{R}^{H\times W\times C}$ denote a clean image and $\mathbf{x}_{t}$ its noisy version at timestep $t\in\{1,\dots,T\}$. The forward diffusion process ($q(\cdot)$) adds Gaussian noise ($\mathcal{N}(\cdot)$) in a Markov chain:
@@
where $\mathbf{c}$ is a conditioning vector, often a CLIP-encoded text prompt. Training minimizes the error between predicted and true noise or the original image $\mathbf{x}_{0}$, depending on parameterization (mean: $\boldsymbol{\mu}_{\theta}$, and covariance $\boldsymbol{\Sigma}_{\theta}$). Latent diffusion models (LDMs) (Wu and De la Torre, 2023; Nam et al., 2024; Zhang et al., 2023b) perform this process in a compressed latent space $\mathcal{Z}$, using an encoder $\mathcal{E}:\mathbb{R}^{H\times W\times C}\rightarrow\mathcal{Z}$ and decoder $\mathcal{D}:\mathcal{Z}\rightarrow\mathbb{R}^{H\times W\times C}$, improving computational efficiency while preserving generation quality.
Style Transfer: Style transfer (Wang et al., 2023a; Hamazaspyan and Navasardyan, 2023; Everaert et al., 2023) aims to modify the appearance of a generated image $\mathbf{x}$ to match the style of a reference image $\mathbf{s}$, while maintaining the content from a source image $\mathbf{c}$. In feature space $\phi(\cdot)$, the style and content losses are defined as:
@@
where $G^{(l)}$ denotes the Gram matrix Li et al. (2017) of features at layer $l$, capturing second-order statistics. The total loss combines both objectives:
@@
Pose Customization with LoRA: Generation of characterization conditions on a specified human pose $\mathbf{p}$, often represented as a 2D keypoint skeleton or heatmap (Wang et al., 2024b; Ju et al., 2023; Khandelwal, 2024). This signal can be embedded and injected into a diffusion model via attention layers or early-stage concatenation. LoRA (Hu et al., 2022) offers an efficient means of adapting large diffusion models to new tasks (e.g., pose-guided generation) by injecting low-rank updates into pretrained weight matrices. For a weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA introduces trainable low-rank matrices $A\in\mathbb{R}^{d\times r}$, $B\in\mathbb{R}^{r\times k}$, modifying the weights as:
@@
where $\alpha$ is a scaling factor and $r\ll\min(d,k)$. By freezing base weights and only training $A$ and $B$, LoRA allows efficient fine-tuning with significantly fewer parameters. LoRA can be applied to attention matrices $W_{q},W_{k},W_{v}$, enabling fast pose-driven customization without full model retraining. In CraftGraffiti instead of using keypoints or heatmap, we extend the input prompt via T5 (Narang et al., 2021) and get the cross-attention map as a guided signal ($B$) for the pose (see Fig.˜2). We maintain the identity of the character through the face-consistent self-attention ($A$) and update the self and cross attention of the diffusion transformer via fine-tuning with LoRA for multi-scale latent feature generation. This multi-scale latent feature is further denoised via VAE in order to get the final edited image in graffiti style with an accurate musical pose described in the text prompt.
@@ 3.1 Why should we perform graffiti style transfer before facial attribute preservation
Theorem. Let $I$ be the input face with attributes $a=F_{a}(I)$. Let $S$ be a diffusion‐style operator and $P$ an operator satisfying $F_{a}(P(X))=a$ for all $X$ (so $P$ restores $I$’s attributes). Then the image $X^{\prime}=P(S(I))$ obeys $F_{a}(X^{\prime})=a$ (it exactly preserves the original attributes), whereas $X^{\prime\prime}=S(P(I))$ satisfies $F_{a}(X^{\prime\prime})=F_{a}(S(I))$ and in general $F_{a}(S(I))\neq a$ unless $S$ itself preserved $F_{a}$. Hence $(P\circ S)(I)$ maintains the facial attributes exactly, whereas $(S\circ P)(I)$ need not.
@@ 3.2 How the extra dimension in the self-attention helps in facial attribute preservation
In Diffusion models (Zhou et al., 2024; Wang et al., 2024c), each self-attention layer computes attention only among features of the current latent (no cross-latent computation). Concretely, the latent feature map of size $H\times W$ is reshaped into a sequence of tokens, and for each token, a query ($Q$), key ($K$), and value ($V$) vector are computed from the same latent features. The attention weights are then obtained by taking dot products of $Q$ and $K$ across all positions (scaled by affinity distance $\sqrt{d}$) and passing through softmax, and the result reweights the $V$ vectors (see Fig.˜3). This means every pixel in the latent try to see every other pixel in the same latent when computing self-attention.
@@ 4 Experimental Results
Implementation Details: We implement CraftGraffiti in PyTorch Paszke et al. (2019), utilizing a pre-trained .1 dev  [^3] https://huggingface.co/spaces/black-forest-labs/FLUX.1-dev (12 billion parameter) transformer specifically designed for graffiti style transfer. The guidance scale is set to 7.5. For preserving facial attributes while fine-tuning the transformer with low rank adaptation, the subject guidance factor ($\lambda$) is set to 0.95, and the style intensity factor is set to 0.7. We set the rank ($\sigma$) to 64 and the regularization factor ($\alpha$) to 128 while fine-tuning the transformer for pose customization. We run the denoising process with 100 steps by default. We only perform latent composition in the first quarter of the denoising process (the first 25 steps). All experiments ran on a single NVIDIA A6000 48GB GPU.
Dataset and Evaluation Metric: For graffiti style generation, we have used the images of the 17K-Graffiti dataset  [^4] https://github.com/visual-ds/17K-Graffiti to fine-tune our style fusion diffusion with LoRA and use it as a pre-trained model during inference. Also, to further validate CraftGraffiti, we have gathered images of people from our laboratory, the Computer Vision Center, Barcelona  [^5] https://www.cvc.uab.es/, with their consent, and use them in the paper for validation and pose customization. Furthermore, to perform a fair qualitative analysis, we test CraftGraffiti and other state-of-the-art approaches with facial feature consistency (FFC) (implementation in the Appendix), aesthetic score (Aes) (Murray et al., 2012) calculated using LAION aesthetic classifier, and Human Preference Score (HPS) (Wu et al., 2023). Also, we have compared them against the inference time to understand their real-time use case scenario.
Qualitative Analysis: We compare CraftGraffiti against several FLUX-based approaches, FLUX + IP-Adapter (Ye et al., 2023), and FLUX.1 Kontext (Labs et al., 2025), however, both approaches fail to generate the graffiti style. Although FLUX + IP-Adapter (Ye et al., 2023) tried to maintain facial consistency, it was unable to change style. On the other hand, FLUX.1 Kontext (Labs et al., 2025) tries to change the background into a more cartoonist style, and the faces are not consistent. This shows how hard to unify the style transfer and consistency tasks and justify our decision to perform style transfer before facial consistency (see Fig.˜4).
@@
We also compare CraftGraffiti with InstructPix2Pix (Brooks et al., 2023); however, it neither adds objects nor generates high-quality images. Similarly, VLMs (Grok 3 (Grok, 3) and GPT4o (Hurst et al., 2024)) maintain consistency, unable to blend style due to the complexity of graffiti style transfer. This empirically shows the complexity of the problem (facial consistency in graffiti style) we are tackling and demonstrates the correctness of the component choice of CraftGraffitii for accurate style blending by preserving the facial features (For more qualitative examples, please refer to the Fig.˜10 of the appendix.)
Quantitative Analysis: Table˜1 presents a quantitative comparison against the state-of-the-art image editing techniques. CraftGraffiti stands in the fourth position after GPT4 (Hurst et al., 2024), GPT4o (Hurst et al., 2024), and (Ye et al., 2023) respectively, in terms of facial feature consistency (FFC), as it is very difficult to compare the preservation of facial features, as most of the feature extraction models (InceptionResNetV1) consider global features rather than the local ones. However, during style transfer, it is hard to preserve the global features of the input images. As other image editing (Grok, 3; Hurst et al., 2024; Ye et al., 2023) models are unable to blend the graffiti style, preserve the global features, and perform better in this metric.
@@
FLUX + IP-Adapter (Ye et al., 2023) 0.8324 3.2414 0.3012 8.2
FLUX.1 Kontext (Labs et al., 2025) 0.6741 2.1749 0.2911 6.7
InstructPix2Pix (Brooks et al., 2023) 0.7112 2.7272 0.1918 3.1
GPT4o (Hurst et al., 2024) 0.8761 4.5193 0.3412 13.4
Grok 3 (Grok, 3) 0.8513 4.1652 0.3102 11.2
@@
On the other hand, Aesthetic Score (Murray et al., 2012; Banerjee et al., 2025) and Human Preference Score (Wu et al., 2023) measure how perfectly the models blend the graffiti style with the input images. CraftGraffiti outperforms all the state-of-the-art image editing techniques in both metrics, as they are unable to blend the graffiti style properly over the input images. Last but not least, CraftGraffiti has very little inference time compared to the VLMs (Grok, 3; Hurst et al., 2024), which makes it suitable for real-time applications.
@@
Similarly, we perform an ablation study on the self-attention (see Fig.˜6) to preserve the facial attribute during pose customization to understand why the extension of the attention dimension helps to preserve the facial features. In order to do that, we took the generated images with CraftGraffiti and pass them through the self-attention layer of FLUX.1 dev, Subject-driven self-attention layer of Consistory (Tewel et al., 2024), and our face consistent self-attention layer. It has been observed that the traditional self-attention of FLUX.1 dev focuses on a certain point, whereas Subject-driven Self-Attention of ConsiStory (Tewel et al., 2024) attends the global features rather than the local ones. On the other hand, our proposed face-consistent self-attention, focused on the local features of the facial attributes, helps to maintain facial attributes during face customization.
Human Evaluation: We have conducted the perceptual study in order to evaluate the credibility and widespread adoption of the graffiti poster synthesized with CraftGraffiti. We chose 60 graffiti posters (20 for each pose) synthesized by (Ye et al., 2023; Labs et al., 2025; Brooks et al., 2023; Grok, 3; Hurst et al., 2024) and CraftGraffiti and asked the users to rate the graffiti on a scale of 1 to 5 (1 Low; 5 high) based on aesthetics, style blending, and recognizability. 47 anonymous users participated in this study. The outcome of this user study is reported in Fig.˜7.